# Prediction of causal genes at GWAS loci with pleiotropic gene regulatory effects using sets of correlated instrumental variables

### v1

@@ Abstract
Here we used causal inference theory to show that MVMR with correlated instruments satisfies the instrumental set condition. This is a classical result by Brito and Pearl (2002) for structural equation models that guarantees the identifiability of individual causal effects in situations where multiple exposures collectively, but not individually, separate a set of instrumental variables from an outcome variable. Extensive simulations confirmed the validity and usefulness of these theoretical results even at modest sample sizes ($n\gtrsim 500$ $1000$). Importantly, the causal effect estimates remain unbiased and their variance small when instruments are highly correlated.
@@ Introduction
More recently, MR has been applied to estimate causal effects of molecular traits such as mRNA or protein abundances on phenotypic traits [7, 8, 9, 10]. While incurring a possible multiple testing burden due to the high number of mRNA or protein measurements that can be made in a single omics experiment, using molecular exposure traits in MR offers potential advantages. This is due to genetic effects on molecular abundances being much larger than those for phenotypic traits, with greater than two-fold allelic effect sizes [11] on gene expression not being uncommon [12].
@@
Moreover, fundamental molecular biological principles strongly limit the potential for horizontal pleiotropy. Indeed, most genetic variants associated to mRNA and protein abundances (and phenotypic traits) are located in non-coding regions of the genome and there are no known mechanisms by which such variants could affect outcome traits other than through affecting transcription and translation of nearby genes, typically considered to be at most a distance of 1 Mbp away [13]. Hence, by limiting the selection of genetic instruments to those in the genomic vicinity of the gene whose mRNA or protein product is used as the exposure in an MR analysis, the risk of horizontal pleiotropy can be reduced significantly.
@@
However, analyses of large transcriptomic datasets have shown that regulatory pleiotropy, where the same genetic variants are associated to multiple nearby genes, is common. For instance, it has been found that around 10% of non-redundant cis-expression quantitative trait loci (cis-eQTLs; single nucleotide polymorphisms (SNPs) associated with the expression of genes within the same genomic locus) were linked to the expression of multiple nearby genes [14]. This covered cases where a true cis-regulatory DNA region is shared between genes as well as cases where genes have distinct regulatory elements that are in high linkage disequilibrium (LD) with one another. More recently, analyses by the GTEx Consortium of RNA abundances across 49 human tissues showed that a median of 57% of variants per tissue were associated with more than one gene in the same locus, typically co-occurring across tissues [12].
@@
Hence, to account for this widespread regulatory pleiotropy, all MR analyses using molecular abundances must employ multivariable MR (MVMR) methods, where all gene products in a genomic locus of interest are treated simultaneously as a set of potential causes for an outcome of interest.
@@
As with single-exposure MR, MVMR was first developed for phenotypic exposure traits [15, 16, 17, 18]. MVMR requires a set of instrumental variable SNPs, at least as many as the number of exposures that are associated with the exposure and outcome variables, and not affecting the outcome other than through the set of exposure variables. MVMR can then estimate the direct causal effect of each exposure on the outcome using a 2-stage least squares method in the single-sample, individual-level data setting, or a “regression of regression coefficients” method in the two-sample summary data setting [15, 16, 17]. Recently, MVMR was applied to transcriptomic and genome-wide association study (GWAS) data in the two-sample setting, in a first attempt to overcome the challenge of regulatory pleiotropy when analyzing mRNA abundances to identify causal disease genes, in a procedure called transcriptome-wide MR (TWMR) [19].
@@
However, the precise conditions to ensure validity of MVMR remain somewhat ambiguous, in particular with regards to the presence of linkage disequilibrium (correlations) between the SNPs in the set of instrumental variables. For instance, Burgess et al.  state that estimates from the 2-stage least squares method in the individual-level data setting are valid even if the genetic variants are in LD [15], but require uncorrelated instruments in their proof of equivalence with the regression method in the summary data setting [16]. McDaid et al.[20] show that the 2-stage least squares estimates can be obtained from univariate summary data even if the genetic variants are in LD, albeit at the cost of no longer being expressed as a “regression of regression coefficients”. Nevertheless, despite their formula for the causal effect estimates being valid for correlated instruments, McDaid et al.[20] still follow common practice in MVMR to select only instruments spread far apart on the genome to eliminate any LD between them. Similarly, Porcu et al.[10], whose TWMR method is based on the analytic results of McDaid et al.[20], only use variants with low mutual LD ($r^{2}\leq 0.1$) in their analysis of loci with regulatory pleiotropy. Moreover, all simulations reported in the literature to test MVMR methods in a controlled situation use independently sampled instruments [15, 16, 17, 19].
@@
To support our theoretical results, we conducted extensive simulations with correlated instruments mimicking real LD from the human genome. We further applied our MVMR method with correlated instrumental variable sets to transcriptomics data from seven vascular and metabolic tissues in 600 coronary artery disease (CAD) cases undergoing surgery from the STARNET study [26] to predict causal genes and tissues at 19 genome-wide significant CAD risk loci with regulatory pleiotropy.
@@
in our SEM. Brito and Pearl [24] derived a sufficient graphical condition for identifiying the paramaters $c_{1},\dots,c_{K}$ by application of Wright’s rule (1), of which we present a simplified version. The set of variables $\{E_{1},\dots,E_{K}\}$ is said to be an “instrumental set” relative to $\{X_{1},\dots,X_{K}\}$ if there exist pairs $(E_{1},p_{1}),\dots,(E_{K},p_{K})$ such that for each $i$ the following three conditions hold.
@@
If $\{E_{1},\dots,E_{K}\}$ is an instrumental set relative to $\{X_{1},\dots,X_{K}\}$, then the parameters of the edges $X_{1}\to Y$, …, $X_{K}\to Y$ are identifiable and can be computed by a set of linear equations [24].
@@
To estimate the causal effects from observational data, we consider multiple finite-sample approximations to the above equations. For finite-sample estimation, the number of instruments can be greater than the number of exposures, that is, we assume we have $N$ observations of the random variables $\{E_{1}.\dots,E_{L}\}$, $\{X_{1}.\dots,X_{K}\}$, and $Y$, with $L\geq K$. We use lower-case letters to denote sample values, e.g. $e_{il}$ denotes the value of $E_{l}$ in observation $i$. We use the notation $e_{i\cdot}$, $e_{.l}$, and $e_{\cdot\cdot}$ to represent the vector of sample values for all $L$ instruments in observation $i$, the vector of all $N$ observations of instrument $l$, and the $N\times L$ matrix of all observations for all instruments, respectively, and similarly for $x_{i\cdot}$, $x_{\cdot k}$, and $x_{\cdot\cdot}$. We assume the observations for all variables are standardized to mean zero and unit standard deviation.
@@
If the model is overdetermined, $L>K$, then in general there is no unique solution to (6), because not all $L$ sample moments will hold exactly. While the least-squares estimator provides one approximate solution to the overdetermined system, [25] proposed an alternative solution, based on bringing the sample moments as close to zero as possible by minimizing the quadratic form
@@
We simulated data from idealized models corresponding to the causal diagrams in Figure 1 where all variables were assumed to be normally distributed.
For the diagrams with a single exposure (Fig. 1A,D), we set the true causal effect $c_{X}=0.3$. In the case of one instrument (Fig. 1A) we varied the instrument strength $a$ between $0.1$ (weak instrument), $0.5$ (mediocre instrument), and $0.8$ (strong instrument). In the case of multiple instruments (Fig. 1D), we simulated ten instruments with randomly sampled instrument strengths, either all strong (effect sizes greater than $0.4$) or all weak (effect sizes smaller than $0.15$).
@@
To test the effect of instrument correlation we simulated data with two instruments (Fig. 1B) where the pairwise correlation of the instruments varied between $0.01$, $0.3$, $0.7$ and $0.9$. The matrix $A=[a_{ij}]$ of instrument effect sizes, where $a_{ij}$ is the direct causal effect of $E_{i}$ on $X_{j}$, was generated randomly such that its determinant was greater than $0.5$ and there were no weak instruments ($a_{ij}$ greater than 0.3).
To test the effect of weak instruments, we simulated data with two instruments (Fig. 1B) with randomly generated non-zero pairwise correlation of the instruments and randomly generated matrix of instrument effect sizes $A$ with determinant either less than $0.2$ or greater than $0.5$.
@@
For more realistic simulations, we simulated data with continuous normally distributed genotype, exposure and outcome variables in such a way that the summary statistics from the simulated data match summary statistics at real GWAS loci for coronary artery disease (CAD). We used eQTL summary statistics from GTEx [12] and CAD GWAS summary summary statistics from the CARDIoGRAMplusC4D meta-analysis [29].
For a given locus of interest with cis-eQTLs $E_{1},\dots,E_{L}$ for genes $X_{1},\dots,X_{K}$, with LD matrix ${\Sigma_{E}}$ (size $L\times L$), matrix of eQTL summary statistics $A$ (size $L\times K$), and vector of CAD GWAS summary statistics $\Sigma_{EY}$ (size $L\times 1$), we simulated data as follows.
For the instruments, we sampled data from the multivariate normal distribution
$E\sim\mathcal{N}(0,{\Sigma_{E}}),$
For the instrument effect sizes, we made the simplifying assumption that all eQTLs are causal for all genes, and set the vector $\mathbf{a}$ of instrument effect sizes to
$\mathbf{a}=(\Sigma_{EE}^{T}\Sigma_{EE})^{-1}\Sigma_{EE}^{T}\sigma_{EX},$ (11)
and sample $X_{1},\dots,X_{K}$ from the conditional distribution
@@
We note that sampling genotype values as binomial variables (representing the number of alternative alleles) is not practical in our situation, because the variables must be correlated according to a predefined LD matrix and the number of such variables ($L$) can be large.
We used the procedure above to create the following simulated datasets.
For the diagram with multiple instruments and a single exposure (Fig. 1D), we considered gene expression of MRPS6 in blood, which has $L=10$ significant eQTLs in GTEx and is the only gene ($K=1$) with eQTLs in this locus (centred on chr 15:35,602,268). The causal effect $c$, in this case, had the value $c=0.64$.
For the diagrams with multiple exposures and correlated instruments (Fig. 1B,E), we used gene expression of NBEAL and FAM117B ($K=2$) in Aorta as exposures, which share $52$ eQTLs, and no other eQTL associations are present in this locus (centred on chr 2: 203029276) in this tissue. Since the determinant of the LD-matrix was very small, we pruned for the most significant CAD instruments by ordering the SNPs based on their significance in the GWAS study and iteratively removing SNPs with lower significance which were in high LD with the more significant SNPs until a minimum threshold for the determinant of the covariance matrix was reached. We compared the situations where we keep only $L=2$ (LD threshold of 0.6), $L=7$ (LD threshold of 0.89), and $L=15$ instruments (LD threshold of 0.96). Causal effects of the genes on the outcome variable $Y$ were calculated to be $c_{NBEAL}=0.24$ and $c_{FAM117B}=-0.08$ when seven instruments were used, and the same values were also used in the other simulations.
For robustness analysis, we repeated the same simulation procedure using gene expression of MRAS and ESYT3 in the Aorta tissue of STARNET as exposures, which share $L=5$ eQTLs in their locus (centred on chr 3:138121920), with simulated causal causal effects $c_{MRAS}=0.208$ and $c_{ESYT3}=-0.294$ on the outcome variable $Y$.
@@
We compared the least squares (eq. (7)) and GMM (eq. (10)) estimators to two other multi-variable MR methods: Transcriptome-Wide MR (TWMR) [19] and the MVMR package [30].
@@
We used eQTL summary statistics from the Stockholm-Tartu Atherosclerosis Reverse Network Engineering Task (STARNET) study, a genetics-of-gene expression study of tissue samples from blood, atherosclerotic-lesion-free internal mammary artery (MAM), atherosclerotic aortic root (AOR), subcutaneous fat (SF), visceral abdominal fat (VAF), skeletal muscle (SKLM), and liver (LIV) obtained during open thorax surgery of 600 coronary artery disease (CAD) patients [26]. eQTL summary statistics from matching tissues in GTEx [12] were used for replication analyses.
We used coronary artery disease (CAD) as the outcome trait and used summary statistics from the CARDIoGRAMplusC4D genome-wide association meta-analysis (GWAMA) study [29]. Summary statistics specific to the European population were extracted using the TwoSampleMR package [31] (study IDs ebi-a-GCST003116).
@@
We iterated through these eQTLs, sorted by GWAS p-value, such that for the first iteration, the first eQTL would be the first lead SNP. We collected all eQTLs within a $0.5Mb$ radius and with LD greater than 0.0 with the lead SNP, and then removed the lead SNP as well as its collection of linked eQTLs from the list, repeating the procedure with the next lead SNP.

### v2

@@ Abstract
Here we used causal inference theory to show that MVMR with correlated instruments satisfies the instrumental set condition. This is a classical result by Brito and Pearl (2002) for structural equation models that guarantees the identifiability of individual causal effects in situations where multiple exposures collectively, but not individually, separate a set of instrumental variables from an outcome variable. Extensive simulations confirmed the validity and usefulness of these theoretical results. Importantly, the causal effect estimates remained unbiased and their variance small even when instruments are highly correlated, while bias introduced by horizontal pleiotropy or LD matrix sampling error was comparable to standard MR.
@@
Author summary
We have conducted simulations to present robust arguments for including instruments in high linkage disequilibrium (LD) in MVMR analyses, especially when independent instruments are unavailable. Although estimates under these conditions exhibit higher variance, they remain unbiased. Our analysis of the principal components of the LD matrix reveals that a significant degree of variation is contributed by SNPs, even at high LD levels which can be exploited for estimation in MVMR settings.
@@ Introduction
More recently, MR has been applied to estimate causal effects of molecular traits such as mRNA or protein abundances on phenotypic traits [7, 8, 9, 10]. Using molecular exposure traits in MR offers potential advantages due to genetic effects on molecular abundances being much larger than those for phenotypic traits, with greater than two-fold allelic effect sizes [11] on gene expression not being uncommon [12]. Moreover, fundamental molecular biological principles reduce the risk of horizontal pleiotropy. Indeed, most genetic variants associated to mRNA and protein abundances (and phenotypic traits) are located in non-coding regions of the genome and there are no known mechanisms by which such variants could affect outcome traits other than through affecting transcription and translation of nearby genes, typically considered to be at most a distance of 1 Mbp away [13]. Hence, by limiting the selection of genetic instruments to those in the genomic vicinity of the gene whose mRNA or protein product is used as the exposure in an MR analysis, the risk of these variants affecting the outcome through alternative pathways not including the exposure gene or protein is reduced.
@@
However, analyses of large transcriptomic datasets have shown that genetic variants with “regulatory pleiotropy”, defined as variants associated with more than one gene in the same locus [12], are common. For instance, it has been found that around 10% of non-redundant cis-expression quantitative trait loci (cis-eQTLs; single nucleotide polymorphisms (SNPs) exhibit regulatory pleiotropy [14]. This covered cases where a true cis-regulatory DNA region was shared between genes as well as cases where genes had distinct regulatory elements that were in high linkage disequilibrium (LD) with one another. More recently, analyses by the GTEx Consortium of RNA abundances across 49 human tissues showed that a median of 57% of variants per tissue are associated with more than one gene in the same locus, typically co-occurring across tissues [12].
@@
When genes or proteins are tested as exposures one-by-one using MR, as in most studies to date, regulatory pleiotropy could clearly lead to horizontal pleiotropy as one cannot know a priori to which of the associated genes a genetic instrument exerts its influence on the outcome. Hence, to account for regulatory pleiotropy, all MR analyses using molecular abundances must employ multivariable MR (MVMR) methods, where all gene products in a genomic locus of interest are treated simultaneously as a set of potential causes for an outcome of interest.
@@
As with single-exposure MR, MVMR was first developed for phenotypic exposure traits [15, 16, 17, 18]. MVMR requires a set of instrumental variable SNPs, at least as many as the number of exposures that are associated with the exposure and outcome variables, and not affecting the outcome other than through the set of exposure variables. MVMR can then estimate the direct causal effect of each exposure on the outcome using a two-stage least squares method in the single-sample, individual-level data setting, or a “regression of regression coefficients” method in the two-sample summary data setting [15, 16, 17]. Recently, MVMR was applied to transcriptomic and genome-wide association study (GWAS) data in the two-sample setting, in a first attempt to overcome the challenge of regulatory pleiotropy when analyzing mRNA abundances to identify causal disease genes, in a procedure called transcriptome-wide MR (TWMR) [19].
@@
However, the precise conditions to ensure validity of MVMR remain somewhat ambiguous, in particular with regards to the presence of linkage disequilibrium (correlations) between the SNPs in the set of instrumental variables. For instance, Burgess et al.  [16] state that estimates from the two-stage least squares method in the individual-level data setting are valid even if the genetic variants are in LD [15], but require uncorrelated instruments in their proof of equivalence with the regression method in the summary data setting. McDaid et al.[20] show that the two-stage least squares estimates can be obtained from univariate summary data even if the genetic variants are in LD, allowing for two-sample MVMR analysis, albeit at the cost of the estimates no longer being expressed as a “regression of regression coefficients”. Nevertheless, despite their formula for the causal effect estimates being valid for correlated instruments, McDaid et al.[20] still follow common practice in MVMR to select only instruments spread far apart on the genome to eliminate any LD between them. Similarly, Porcu et al.[19], whose TWMR method is based on the analytic results of McDaid et al.[20], only use variants with low mutual LD ($r^{2}\leq 0.1$) in their analysis of loci with regulatory pleiotropy. Moreover, all simulations reported in the literature to test MVMR methods in a controlled situation use independently sampled instruments [15, 16, 17, 19].
@@
To support our theoretical results, we conducted extensive simulations with binomially distributed, correlated instruments mimicking real LD from the human genome. We further applied our MVMR method with correlated instrumental variable sets to transcriptomics data from seven vascular and metabolic tissues in 600 coronary artery disease (CAD) cases undergoing surgery from the STARNET study [26] to predict causal genes and tissues at 19 genome-wide significant CAD risk loci with regulatory pleiotropy.
@@
Fig 1: Causal diagrams for identification (A-C) and estimation (D-F) of causal effects at GWAS loci with regulatory pleiotropy. (A) The standard instrumental variable graph is only suitable for modelling the causal effect of gene $X$ on outcome trait $Y$ if $X$ is the only cis-eGene in the locus of the genome-wide significant variant $E$ for $Y$. (B) A causal graph for MVMR with regulatory pleiotropy, where two or more cis-eGenes $X_{i}$ are found in a genome-wide significant locus for $Y$; the causal effects of each $X_{i}$ on $Y$ can be identified if there exist at least the same number of causal variants as the number of cis-eGenes in the GWAS locus, even when they are in LD with each other, because the variants form an instrumental set. (C) If the number of causal variants is smaller than the number of cis-eGenes and other variants in the locus are merely associated by LD, the causal effects of $X_{i}$ on $Y$ are non-identifiable. (D,E) When estimating causal effects from finite samples, we use an overdetermined set of genetic variants as instruments, without knowing the number or identity of the causal variants. (F) If the (unknown) number of causal variants in a locus is smaller than the number of cis-eGenes, causal effects can not be estimated, a condition that can be tested by analyzing the rank or determinant of the covariance matrix between the variants and cis-eGenes. Solid arrows represent direct causal effects and dashed (bi)directed arrows represent non-zero covariances due to unobserved ($U$) confounders.
@@
in our SEM. Brito and Pearl [24] derived a sufficient graphical condition for identifying the parameters $c_{1},\dots,c_{K}$ by application of Wright’s rule (1), of which we present a simplified version. The set of variables $\{E_{1},\dots,E_{K}\}$ is said to be an “instrumental set” relative to $\{X_{1},\dots,X_{K}\}$ if there exist pairs $(E_{1},p_{1}),\dots,(E_{K},p_{K})$ such that for each $i$ the following three conditions hold.
@@
If $\{E_{1},\dots,E_{K}\}$ is an instrumental set relative to $\{X_{1},\dots,X_{K}\}$, then the parameters of the edges $X_{1}\to Y$, …, $X_{K}\to Y$ are identifiable and can be computed by a set of linear equations [24]. For concrete examples verifying the instrumental set conditions, see the Supplementary Information.
@@
To estimate the causal effects from observational data, we consider multiple finite-sample approximations to the above equations. For finite-sample estimation, the number of instruments can be greater than the number of exposures, that is, we assume we have $N$ observations of the random variables $\{E_{1},\dots,E_{L}\}$, $\{X_{1},\dots,X_{K}\}$, and $Y$, with $L\geq K$. We use lower-case letters to denote sample values, e.g. $e_{il}$ denotes the value of $E_{l}$ in observation $i$. We use the notation $e_{i\cdot}$, $e_{.l}$, and $e_{\cdot\cdot}$ to represent the vector of sample values for all $L$ instruments in observation $i$, the vector of all $N$ observations of instrument $l$, and the $N\times L$ matrix of all observations for all instruments, respectively, and similarly for $x_{i\cdot}$, $x_{\cdot k}$, and $x_{\cdot\cdot}$. We assume the observations for all variables are standardized to mean zero and unit standard deviation.
@@
If the model is overdetermined, $L>K$, then in general there is no unique solution to (6), because not all $L$ sample moments will hold exactly. While the least-squares estimator provides one approximate solution to the overdetermined system, Hansen [25] proposed an alternative solution, based on bringing the sample moments as close to zero as possible by minimizing the quadratic form
@@
We simulated data from idealized models corresponding to the causal diagrams in Figure 1 with binomially distributed instruments and normally distributed exposure and outcome variables.
For the diagrams with a single exposure (Fig. 1A,D), we set the true causal effect $c_{X}=0.3$. In the case of one instrument (Fig. 1A) we varied the instrument strength $a$ between $0.1$ (weak instrument), $0.5$ (mediocre instrument), and $0.8$ (strong instrument). In the case of multiple instruments (Fig. 1D), we simulated ten instruments with randomly sampled instrument strengths, either all strong (effect sizes greater than $0.5$) or all weak (effect sizes smaller than $0.03$).
@@
To test the effect of instrument correlation we simulated data with two instruments (Fig. 1B) where the pairwise correlation of the instruments varied between $0.01$, $0.3$, $0.7$, $0.9$ and $0.95$. The matrix $A=[a_{ij}]$ of instrument effect sizes, where $a_{ij}$ is the direct causal effect of $E_{i}$ on $X_{j}$, was generated randomly such that its determinant was greater than $0.5$ and there were no weak instruments ($a_{ij}$ greater than 0.6).
To test the effect of weak instruments, we simulated data with two instruments (Fig. 1B) with randomly generated non-zero pairwise correlation of the instruments and randomly generated matrix of instrument effect sizes $A$ with determinant either less than $0.2$ and effect size of each instrument less than $0.03$ or determinant greater than $0.5$ and effect size of each instrument greater than $0.5$.
@@
For more realistic simulations, we simulated data with binomially distributed genotypes and normally distributed exposure and outcome variables in such a way that the summary statistics from the simulated data match summary statistics at real GWAS loci for coronary artery disease (CAD). We used eQTL summary statistics from GTEx [12] and CAD GWAS summary summary statistics from the CARDIoGRAMplusC4D meta-analysis [29].
For a given locus of interest with cis-eQTLs $E_{1},\dots,E_{L}$ for genes $X_{1},\dots,X_{K}$ ($L\geq K$), with LD matrix ${\Sigma_{E}}$ (size $L\times L$), matrix of eQTL summary statistics $A$ (size $L\times K$), and vector of CAD GWAS summary statistics $\Sigma_{EY}$ (size $L\times 1$), we simulated data as follows.
To sample binomially distributed genetic instruments according to a predefined LD matrix, we made the simplifying assumption that the correlation structure between the simulated instruments can be described by a Markov chain. We sampled the first SNP’s genotype from a binomial distribution with probability of success equal to its real minor allele frequency (MAF). For every other SNP, we sampled a genotype from a binomial distribution with probability of success conditional on the genotype of the previous SNP in the sequence, such that both the MAF and pairwise LD between successive SNPs match the real data from the simulated locus. Further details are given in the Supplementary Information.
We assumed that a randomly selected subset of at least $K$ eQTLs were causal and all others were associated with the exposures only through LD. For the causal eQTLs, we sampled the vector $\mathbf{a}$ of instrument effect sizes randomly between $0.7$ and $1.0$,
@@
We performed simulations to analyze the influence of LD matrix accuracy, weak instrument bias, standard error calculation method, and horizontal pleiotropy.
For the simulations on the accuracy of the LD matrix, we generated datasets for the SLC22A3-LPA-PLG locus (centered on chr $6:161089307$) in liver as exposures. Causal effects of the genes on the outcome variable $Y$ were calculated to be $c_{\textit{SLC22A3}}=0.15$, $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ when seven instruments were used, and the same values were also used in the other simulations. We compared the situations where $L=3$ (LD threshold of 0.25), $L=4$ (LD threshold of 0.3), $L=5$ (LD threshold of 0.4), $L=6$ (LD threshold of 0.5), and $L=7$ instruments (LD threshold of 0.8) were used. We performed one-sample simulations where the sample size varied between $500-30000$. For the instrument effect sizes, we assumed that three eQTLs were causal for all genes (their effect sizes sampled from a uniform distribution within the range $0.7-1.0$), while additional eQTLs were assumed to be associated with the exposures only through LD.
For the simulations on weak instrument bias, we generated datasets within the same locus and setup as the LD accuracy simulations except here we kept $L=8$ eQTLs (LD threshold of 0.96) and varied the sample size between $500-10000$. For the instrument effect sizes, we again assumed that only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.7-1.0$ for strong instruments and within the range $0.001-0.03$ for weak instruments. We confirmed that these instruments were indeed weak/strong by employing the conditional F-statistics computed using the Mendelian Randomization package [30] as well as the MVMR package [31] (S3 Fig).
For the simulations on comparing the standard errors computed approximately using summary-level data and exactly using individual-level data, we generated datasets within the same locus and setup as the previous simulations. We kept $L=7$ eQTLs (LD threshold of 0.95) and varied the sample size between $500-10000$. For the instrument effect sizes, we again assumed that only three eQTLs were causal for all genes. Details about the standard error calculation using summary and individual-level data are given in detail in the Supplementary Information.
To simulate the effect of horizontal pleiotropy, we generated datasets for the SLC22A3-LPA-PLG locus with causal effects of the genes on the outcome variable $Y$ set to $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ and $c_{\textit{SLC22A3}}\in[0.0,0.4]$. We used eight instruments where only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.7-1.0$. We estimated the effects in a misspecified estimation model where we performed MVMR only using LPA-PLG genes, assuming horizontal pleiotropy through one unknown/unmeasured gene SLC22A3 in the same locus.
To simulate the effect of potential bias introduced by two-sample MR, we generated independently sampled eQTL and GWAS datasets for the MRAS-ESYT3 locus in the Aorta tissue of STARNET, which share $L=5$ eQTLs in their locus (centred on chr 3:138121920), with causal effects of the genes on the outcome variable $Y$ set to $c_{\textit{MRAS}}=0.208$ and $c_{\textit{ESYT3}}=-0.294$. We simulated eQTL sample sizes between $300-4000$ and GWAS sample sizes between $10000-140000$.
@@
We compared the least squares (LS) (eq. (7)) and GMM (eq. (10)) estimators to three other multi-variable MR methods: Transcriptome-Wide MR (TWMR) [19], the mv_multiple estimator from TwoSampleMR package [32], and the mv_mvivw estimator from the Mendelian Randomization package [30].
@@
We used eQTL summary statistics from the Stockholm-Tartu Atherosclerosis Reverse Network Engineering Task (STARNET) study, a genetics-of-gene expression study of tissue samples from blood ($n=481$), atherosclerotic-lesion-free internal mammary artery (MAM, $n=552$), atherosclerotic aortic root (AOR, $n=539$), subcutaneous fat (SF, $n=573$), visceral abdominal fat (VAF, $n=531$), skeletal muscle (SKLM, $n=534$), and liver (LIV, $n=576$) obtained during open thorax surgery of 600 coronary artery disease (CAD) patients [26]. eQTL summary statistics from matching tissues in GTEx [12] were used for replication analyses.
We used coronary artery disease (CAD) as the outcome trait and used summary statistics from the CARDIoGRAMplusC4D genome-wide association meta-analysis (GWAMA) study [29]. Summary statistics specific to the European population were extracted using the TwoSampleMR package [32] (study IDs ebi-a-GCST003116, $n=141,217$).
@@
We iterated through these eQTLs, sorted by GWAS p-value, such that for the first iteration, the first eQTL would be the first lead SNP. We collected all eQTLs within a $0.5Mb$ radius and with non-zero LD with the lead SNP, and then removed the lead SNP as well as its collection of linked eQTLs from the list, repeating the procedure with the next lead SNP. We computed conditional F-statistics using the Mendelian Randomization [30] and MVMR [31] packages to verify instrument strength.
@@
We computed summary-based standard errors using the sample size from the GWAS study (see Supplementary Information for details) and obtained p-values from its asymptotic normal distribution. We confirmed using simulations that summary-based and individual-level based standard errors were comparable (S4 Fig). We further confirmed standard errors using the “mr_mvivw” method of the “MendelianRandomization” package (Supplementary Figure S6 Fig). We note that to the best of our knowledge, the only standard error estimator that takes into account variability due to the different sample sizes in the two-sample setting [33] requires individual-level data and can thus not be computed in the two-sample summary-data setting.